# FocusFlow: 3D Gaze-Depth Interaction in Virtual Reality Leveraging Active Visual Depth Manipulation

### v2

@@ 10.2. Compatibility and Extendibility
The introduction of visual depth as a unique input dimension does not clash with other input modalities such as eye movements or blinks. Instead, FocusFlow can be seamlessly integrates with other gaze interactions. Moreover, the visual depth modality can be harmoniously coupled with other VR interaction modes like head and hand movements. For instance, a user might activate a virtual window using visual depth transition and subsequently interact with components within the window using hand gestures. Such rich and compound interactions are ripe avenues for exploration in future research.
10.3. Gaze Depth as a Continuous Input
@@
The evaluations in Section 3.3 show that high-accuracy depth estimations are limited to close distances. While this limits the virtual windows to certain depths, we can expand this interaction method to a continuous input by tracking the continuous depth changes within a certain range. This will enable various new inputs such as scrolling on a webpage or rotating a 3D object by looking at micro-gaze depth changes. In addition, different mapping functions can be defined to match the gaze depth changes to the speed of interactions.
10.4. Virtual Window Depth Adaptability
@@
To elevate user experience and reduce erroneous virtual window activation, the position of the interaction layer could be adapted to a target’s location. For example, when the target is a few meters away from the user, the interaction layer could be set to one meter in front of the user. However, if the target distance is less than one meter, the position of the interaction layer will be automatically set to half a meter to the user. However, such adaptability necessitates visual cues to inform users of the adjusted window position, ensuring seamless interactions. In addition, in this paper, we assumed that the virtual window depth is always displaces closer to the user than the interactive object. However, FocusFlow can be seamlessly extended to applications where virtual windows can be designed further from the interactive object. For example, in a surgical training application, the user can switch between observing different layers inside the patients body by actively changing their focal depth further from the upper surface of the body.
@@ Appendix A Autostereogram
Figure 21. Autostereogram Example. When you put your gaze point sightly further and make the green circle and pink circle overlap, you can see a 3D image. Making the image as large as possible and not too close to the image will help you see the 3D image successfully.

### v3

@@ 10. Limitation and Future Work
Multi-modal Interaction Inputs: Visual depth is a unique input dimension with a marginal likelihood of clashing with other eye behaviors. However, it only offers limited interaction inputs such as selection or zoom-in. To improve the number of inputs, FocusFlow can be seamlessly integrated with other eye input methods, such as eyelid movements or blinks. Moreover, the visual depth modality can be harmoniously coupled with other VR interaction modes like hand and head inputs. For instance, a user might activate a virtual window through gaze-depth interaction and subsequently interact with components within the window using direct or indirect hand gestures (Pfeuffer, 2017). Such compound interactions can be explored in detail in future research.
@@
Gaze Depth as a Continuous Input: The evaluations in Section 3.3 show that high-accuracy depth estimations are limited to close distances. While this limits our accurate depth detection over long distances, we can still utilize depth data at close range for more fine-grained inputs. For example, we can expand our gaze-depth interaction from binary selection to continuous inputs by continuously tracking exact visual depth changes and mapping the depth to some operational values. This will enable various manipulations such as scrolling on a webpage (Sharmin et al., 2013) or rotating a 3D object (Liu et al., 2020) by looking at micro gaze-depth changes.
@@
Virtual Window Depth Adaptability: To elevate user experience and reduce erroneous virtual window activation, the position of the interaction layer could be adapted to a target’s location. For example, when the target is a few meters away from the user, the interaction layer could be set to one meter in front of the user. However, if the target distance is less than one meter, the position of the interaction layer will be automatically set to half a meter from the user. However, such adaptability necessitates visual cues to inform users of the adjusted window position, ensuring seamless interactions.
(a) Looking at the object
(b) Looking through the object
Figure 21. Surgical training application.
Looking Through Interaction: In this paper, we assumed that the virtual window is always placed closer to the user than the interactive object. However, FocusFlow can be seamlessly extended to applications where information is designed further from the interactive object. For example, in the surgical training application shown in Figure 21, the user can switch between observing different layers inside the patient’s body by actively changing their focal depth further from the upper surface of the body.
@@ 11. Conclusion
Acknowledgements.
This research was partially funded by UIUC-Insper program. We thank Professor Andrew Kurauchi and Professor Luciano Soares for their guidance in eye-tracking characterization and Arnav Shah for designing and implementing the surgical training application.